OpenStudy (kobeni-chan):

***FAN and MEDAL!*** I need to find a graphing website that will take restrictions such as: x = 2t y = t + 5, **-2 ≤ t ≤ 3** I've tried a couple sites already and I can't figure out how to do it. Fan and medal for someone who can give me a website and tell me how to put those values in, and it work. Please, I need to finish this in the next hour!

OpenStudy (cruffo):

Have you tried wolfrram alpha's parametric plot ? https://www.wolframalpha.com/input/?i=parametric+plot+(2t,+t+%2B+5)+where+-2+%3C%3D+t+%3C%3D+3

OpenStudy (kobeni-chan):

No I haven't! thank you so much!!

OpenStudy (michele_laino):

@kobeni-chan if we solve the first equation for \(t\), we get: \[t = \frac{x}{2}\] now I substitute such value for \(t\) into the second equation, and I can write: \[y = \frac{x}{2} + 5\]

OpenStudy (michele_laino):

so, you have to draw a line whose cartesian equation is: \[y = \frac{x}{2} + 5\]

OpenStudy (michele_laino):

furthermore, since: \[ - 2 \leqslant t \leqslant 3\] after the substitution, we get: \[ - 4 \leqslant x \leqslant 6\]

OpenStudy (michele_laino):

so, please draw a line \(y=x/2 +5\), wherein \[ - 4 \leqslant x \leqslant 6\]